# Knowledge Distillation Driven Semantic NOMA for Image Transmission with Diffusion Model

### v1

@@ V-D Ablation Study
VI Conclusion
This paper proposed the KDD-SemNOMA scheme, a pioneering framework for wireless image semantic NOMA transmission. By integrating adaptive channel-aware encoding, knowledge distillation optimization, and diffusion model-based generative refinement, our approach effectively addressed the critical challenges of semantic feature interference and perceptual quality degradation under bandwidth-constrained multi-user scenarios. Our proposed KDD-SemNOMA leveraged a ConvNeXt-based architecture with enhanced AF-Module to ensure robust performance over AWGN and Rayleigh fading channels. At the first stage, knowledge distillation strategy leveraged interference-free orthogonal transmission teacher model to optimize SemNOMA student model training, yielding PSNR and SSIM gain without additional inference complexity. At the second stage, the diffusion model-based refinement harnessed generative priors to elevate perceptual quality, transforming initial reconstructions into high-fidelity images. Experimental results demonstrate KDD-SemNOMA’s superiority over state-of-the-art baselines, simultaneously advancing pixel-level fidelity and perceptual quality. Future work will explore scalability to larger user populations, real-time implementation under dynamic channel conditions, and integration of multimodal data.

### v2

@@ VI-D Ablation Study
VII Conclusion
This paper proposed the KDD-SemNOMA scheme, a pioneering framework for wireless image semantic NOMA transmission. By integrating adaptive channel-aware encoding, knowledge distillation optimization, and diffusion model-based generative refinement, our approach effectively addressed the critical challenges of semantic feature interference and perceptual quality degradation under bandwidth-constrained multi-user scenarios. Our proposed KDD-SemNOMA leveraged a ConvNeXt-based architecture with enhanced AF-Module to ensure robust performance over AWGN and Rayleigh fading channels. At the first stage, knowledge distillation strategy leveraged interference-free orthogonal transmission teacher model to optimize SemNOMA student model training, yielding PSNR and SSIM gain without additional inference complexity. At the second stage, the diffusion model-based refinement harnessed generative priors to elevate perceptual quality, transforming initial reconstructions into high-fidelity images. Experimental results demonstrate KDD-SemNOMA’s superiority over state-of-the-art baselines, simultaneously advancing pixel-level fidelity and perceptual quality. Future work will extend the scenario from uplink to downlink NOMA, explore scalability to larger user populations, adjust the compression rate based on real-time channel state information, and integration of multimodal data.